### v2

# MACEval: A Multi-Agent Continual Evaluation Network for Large Models

Abstract
Hundreds of benchmarks dedicated to evaluating large models have been presented over the past few years. However, most of them remain closed-ended and are prone to overfitting due to the potential data contamination. Moreover, the increasing scale and scope of current benchmarks with transient metrics, as well as the heavily human-dependent curation procedure, pose significant challenges for timely maintenance and adaptation. In this paper, we introduce MACEval, a Multi-Agent Continual Evaluation network for dynamic evaluation of large models, and define new metrics to quantify performance longitudinally. MACEval employs an interactive and autonomous evaluation mode, utilizing role assignment, in-process data generation, and evaluation routing through a cascaded agent network. Extensive experiments on 23 large models demonstrate the effectiveness of MACEval, which also lightens the evaluation process and reduces a considerable amount of overhead. We hope that MACEval can broaden future directions of large model evaluation. Project page: https://github.com/zijianchen98/MACEval.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) have achieved unprecedented performance in diverse applications such as visual understanding (Achiam et al., 2023; Kavukcuoglu, 2025; Zhu et al., 2025; Chen et al., 2025), complex long-text reasoning (math and coding) (Anthropic, 2024; Grattafiori et al., 2024; Team, 2025), and spatial cognition (Wu et al., 2024b; Yang et al., 2024b). Evaluation holds an equally pivotal role as pre-training in the development of large models, functioning not only as a benchmark for understanding current performance but also as a critical pillar for aligning model capabilities with intended objectives. In this regard, it is crucial to construct transparent, trustworthy, and zero-contamination evaluation processes.
Figure 1: Paradigm comparison between current large model benchmarks and the proposed MACEval.
Despite the proliferation of various evaluation benchmarks, doubts about the genuine capabilities of large models persist (Li et al., 2024c; Yang et al., 2024c), particularly when discrepancies arise between leaderboard scores and real-world user experience. As shown in Fig. 1, we conclude the following three fundamental issues from the perspectives of data source, evaluation pipeline, and evaluation metric: 1) Closed-ended data. Most benchmarks, alongside a corresponding well-organized dataset, which contains evaluation materials (e.g., texts, images, or other modalities) collected from the Internet or existing large-scale datasets, pose potential data contamination risks of overlap with the training set (Xu et al., 2024; Deng et al., 2024). Once constructed, such static datasets are susceptible to obsolescence due to overfitting, especially considering the current rapid monthly update cycle of large models. 2) Human-dependent evaluation pipeline. As the scale and coverage of current benchmarks continue to expand, the drawbacks of heavily human-involved annotation (e.g., query-answer design and data refinement) and result review that traditional benchmarks (Wu et al., 2024a; Liu et al., 2024b; Yue et al., 2024) rely on are gradually being exposed. Its time-consuming and high-cost characteristics hinder timely maintenance and updates to catch up with the models’ evolving performance. 3) Limited and temporary metric. The current performance quantification strategies that are limited by closed-ended datasets may not reflect the model’s maximum capabilities. Besides, many metrics are transient in nature and fail to incrementally reflect performance changes for future models, making longitudinal assessment challenging. Recently, some researchers have proposed to address data-related issues by employing techniques such as vision-language bootstrapping (Yang et al., 2024c), establishing specific data generation rules (Zhu et al., 2024a), or meta probing agent (Zhu et al., 2024b). However, solutions for improving evaluation pipelines and metrics remain largely unexplored.
To address these problems, we introduce MACEval, a dynamic continual evaluation framework that measures the progress of large models autonomously by implementing a multi-agent collaboration system. First, we model the traditional one-way problem-solving evaluation process as an interactive interview process and assign three distinct roles in MACEval, i.e., the interviewee, the interviewer, and the supervisor, who are responsible for answering, developing the questions, and inspecting the whole evaluation process, respectively. In particular, adhering to the design principles of open-ended evaluation tasks, we deploy reliable and contamination-free data sources by integrating an in-process generation strategy with real-world rule-based supervision, thus reducing unnecessary data collection and evaluation expenditures. Furthermore, instead of relying on human procedures, MACEval builds an agent-based evaluation network with a message-passing mechanism to automatically perform assessments across different tasks through pre-defined evaluation topologies, which enables hierarchical capability evaluation and is more flexible. Based on this, we propose an Area Under Curve (AUC)-inspired evaluation routing-based metric to measure the overall performance of a large model continually and maximally. Experiments on 23 popular large models across five typical capabilities demonstrate the effectiveness and efficiency of the proposed MACEval.
Our contributions are summarized in three-fold:
  • We introduce MACEval, a multi-agent continual evaluation network for dynamic large model assessment. By modeling the evaluation procedure as an autonomous interview, MACEval effectively mitigates data contamination and enhances evaluation efficiency.
  • We propose an AUC-inspired metric for large models to evaluate longitudinally, which can be incorporated with the evaluation routing to assess the overall performance utilization rate.
  • We evaluate the proprietary and open-source LLMs and MLLMs on several open-ended tasks to demonstrate the effectiveness of the proposed MACEval. Evaluation results demonstrate the superior efficiency, flexibility, and scalability of MACEval compared to existing benchmarks, providing valuable insights for future research on large model evaluation.
2 Related Work
Data Contamination. Recently, data contamination issues have gained widespread attention across evaluation benchmarks for both LLMs and MLLMs (Carlini et al., 2023; Xu et al., 2024; Li & Flanigan, 2024). Researchers from OpenAI and the Llama group conducted contamination studies for GPT-4 (Achiam et al., 2023) and Llama2 (Touvron et al., 2023) on their pre-training data. Study (Li et al., 2024c) reveals significant contamination rates in academic exam-based benchmarks such as MMLU (29.1%) (Hendrycks et al., 2021) and C-Eval (45.8%) (Huang et al., 2023), primarily attributed to the widespread dissemination and circulation of academic test questions. Deng et al. (Deng et al., 2024) proposed a corpora overlap investigation protocol, TS-Guessing, and detected 57% exact match rate of ChatGPT in predicting masked choices in the MMLU test set. Yang et al. (Yang et al., 2024c) reported image overlaps of over 84.4% and 33.2% between SEEDBench (Li et al., 2024a) and pre-training datasets LAION-100M (Schuhmann et al., 2021) and CC3M (Sharma et al., 2018), respectively. Such exposure can significantly impact the reliability of the evaluation, leading to inflated results that do not accurately reflect the true capabilities of the large models. In this paper, we address this problem by reforming the source of evaluation content, introducing AI-generated data, and an open-ended task design strategy to enhance benchmark robustness.
Figure 2: An overview of our proposed MACEval, which consists of three primary phases: evaluation capability determination, MAEN construction, and open-ended task selection. The pipeline models the evaluation of large models as a multi-round interview process.
Dynamic Evaluation. One recent promising attempt to mitigate the issue of data contamination in large model evaluations is dynamic evaluation. (Wang et al., 2025) implemented several perturbations, such as paraphrasing, adding noise, and reversing polarity, to construct evolving instances for testing LLMs against diverse queries and data noise. DyVal (Zhu et al., 2024a) proposes to dynamically generate evaluation samples under pre-defined constraints, which modulates a graph-based algorithm generation structure and fine-grained control over problem difficulty by adjusting the structural complexity. Afterwards, (Zhu et al., 2024b) presented meta-probing agents, which automatically refresh an original evaluation problem following psychometric theory on three basic cognitive abilities, including language understanding, problem solving, and domain knowledge. In (Yang et al., 2024c), the authors designed various bootstrapping strategies (e.g., image editing and sentence rephrasing) with complexity control for both image and question modification. However, these studies primarily focus on modifying the sources of evaluation data, without achieving fully dynamic and automatic evaluation processes. Moreover, the corresponding evaluation metrics remain static and transient, failing to reflect the dynamic evolution characteristics of large models adequately.
3 The MACEval
3.1 Design Principles
Focusing on Autonomous Evaluation Process. Unlike existing LLM or MLLM benchmarks that are strongly human-involving in source content collection and relatively independent in cross-ability evaluation, our MACEval follows two basic principles: (1) Mitigating the need for pre-curated evaluation sets, all visual and textual query–answer pairs are generated on-the-fly; We conduct a preliminary exploration on 9 tasks across five key domains currently emphasized in the field, i.e., visual perception, textual comprehension, math, algorithms, and coding (White et al., 2025; Zhang et al., 2025). As illustrated in Fig. 3, each task incorporates dynamic difficulty variables (e.g., distortion level, string length, or cyclomatic complexity) to enable quantitative assessment. (2) Progressive capability evaluation scheme with real-time task adjustment. We adhere to the principles in building the MACEval, while making it applicable to evaluating all-round abilities. More details are in App. A.
Pushing to the Limit. Since existing benchmark datasets are finite and closed-ended, the measured performance may not reflect the model’s maximum capabilities. To push the evaluated model to the limit, we adopt a stress-testing strategy, where the model is continuously challenged with increasingly difficult query-answer tasks until it fails to provide a correct response. Henceforth, we can derive a long-standing metric by iteratively updating the envelope area formed by connecting performance points across different difficulty levels.
Figure 3: A data card of 10 open-ended tasks that evaluate the visual perception, text comprehension, math, algorithm reasoning, and coding abilities of large models. The difficulty variable of each task is indicated in red.
3.2 Definition and Formulation
We define the input space of a multi-agent system as $\mathbb{I}=(\mathbb{M},\mathbb{R},\mathbb{T})$, where $\mathbb{M}$ denotes the candidate pool of LLMs or MLLMs. $\mathbb{R}$ denotes the set of pre-defined agent roles (e.g., interviewee, interviewer, and supervisor), and $\mathbb{T}$ is the set of evaluation topologies, i.e. collaboration modes, such as point-to-point link, tree, star, and multi-hop network. Within the evaluation space, a Multi-Agent Evaluation Network (MAEN) is defined as follows:
Definition 1 (Multi-Agent Evaluation Network): The MAEN $\mathcal{G}$ includes several large models with distinct identities (node type $\mathcal{M}$), participating collaboratively (edge type $\mathcal{R}$) in a configurable (topological type $\mathcal{T}$) evaluation process:
$\mathcal{G}=\left\{{\{{\mathcal{M}_{i}}\}_{i=1}^{M},\{{\mathcal{R}_{j}}\}_{j=1}^{R},\{{\mathcal{T}_{k}}\}_{k=1}^{T}}\right\},$ (1)
where $M$, $R$, and $T$ are the number of each corresponding component.
Definition 2 (Multi-Agent Continual Evaluation Stream): The MACEval stream can be denoted by a mapping function $f:\mathbb{M}\times\mathbb{R}\times\mathbb{T}\to\mathcal{G}$ that maps the input space $\mathbb{I}$ to an MAEN $\mathcal{G}$ tailored for the evaluated capability $\mathcal{C}$ from a non-stationary data stream $\mathcal{S}=\left\{{{{({q,a})}_{0}},{{({q,a})}_{1}},\cdots,{{({q,a})}_{\infty}}}\right\}$, where data tuple $({q,a})_{t}$} (query-answer pair) is dynamically generated via a step $t$-dependent function.
[…]
Specifically, when the message passing terminates at a supervisor node, the aggregated messages store the information of the entire MAEN.
3.3 MACEval Framework
As shown in Fig. 2, the proposed MACEval network is configured by three main factors: 1) the role assignment defines the basic functionality of network nodes; 2) the evaluation topology determines the sequential order and interdependencies within the evaluation process; 3) the evaluation network-based metrics and the evaluating continually mechanism provides comprehensive and sustainable capability assessment.
Table 1: Performance comparison on four text-only capabilities under a 1-hop line evaluation topology. We report the ACC-AUC/Maximum performance, as well as the utility-capability ratio. The best results are highlighted in bold.
Interviewee (LLM) Interviewer Text Comprehension Math Algorithm Coding $\epsilon_{\text{overall}}$
STUdisrupt. STUmask. SP Arith.scale Arith.oper. BTT SPSnode SPSedge Gener. Debug.
GPT-4o GPT-4o 8.857/10 6.043/7 2.6/6 1.7/2 0.4/1 3.1/4 5.5/10 4.1/6 6.1/11 14.3/16 6.889
GPT-4.1 8.438/9 4.815/6 3.6/10 1.8/2 0.3/1 2.6/4 5.2/10 3.7/6 6.3/12 14.8/16 6.537
Gemini 1.5 Pro 8.766/10 5.289/6 1.4/3 1.3/2 0.5/1 2.8/3 5.4/10 4.2/6 6.1/12 14.6/18 6.868
Gemini 2.0 Flash 8.898/10 5.049/6 2.8/8 1.6/2 0.6/1 3.1/4 6.1/10 4.5/7 7.3/13 16.6/22 6.825
Gemini 2.5 Pro 8.978/10 6.062/7 3.8/9 3.3/5 0.9/3 3.2/4 6.2/10 4.9/8 7.8/15 17.8/22 6.508
DeepSeek-V3 5.293/6 5.099/6 2.0/5 2.6/4 0.7/1 3.0/3 4.4/10 4.6/6 6.4/13 15.3/17 7.081
DeepSeek-R1 9.871/10 5.936/7 8.3/10 7.2/10 0.9/1 3.1/5 9.3/13 7.2/9 7.5/13 17.9/20 7.892
Qwen3-8B GPT-4o 7.626/10 2.846/4 1.4/4 1.0/1 0.9/2 1.1/2 4.2/8 2.8/4 5.8/8 11.5/14 6.596
Qwen2.5-7B 6.797/8 3.003/4 0.9/2 0.9/1 0.5/1 1.6/2 3.4/8 2.6/4 5.2/8 9.8/13 6.729
Qwen2.5-14B 7.903/10 3.327/4 2.0/5 1.0/1 0.5/1 2.2/3 4.4/8 3.0/4 5.5/8 11.2/13 7.104
Qwen2.5-72B 8.543/10 4.274/5 2.2/5 1.0/1 0.5/1 2.4/3 2.8/4 3.7/5 5.8/9 13.9/15 7.460
Qwen2-7B 7.339/9 3.053/4 0.5/1 0.7/1 0.4/1 1.2/2 2.8/7 2.4/4 4.7/7 8.5/13 6.104
Llama3.3-70B 7.619/10 4.564/6 2.8/8 1.0/1 0.5/1 2.6/3 4.1/8 3.3/4 5.7/9 13.6/14 7.181
Llama3.2-3B 5.641/8 2.369/3 0.1/1 0.3/1 0.3/1 0.6/1 1.4/4 1.7/3 2.2/5 7.3/8 5.064
Llama3.1-8B 6.740/9 3.006/4 0.7/4 0.2/1 0.4/1 0/0 2.9/8 2.1/4 4.6/7 9.3/12 4.595
Role Assignment. We mainly involve three types of agents, i.e., the interviewer, the interviewee, and the supervisor. Among them, the interviewee, namely the evaluated model, generates responses according to the given queries. The interviewer performs real-time performance evaluations and determines whether to transform a given question into a new one with dynamically increasing difficulty based on the interviewee’s responses. To ensure the validity of the evaluation process and to avoid scenarios such as repetitive questioning or posing questions that do not meet evaluation requirements, we deploy a third-party model distinct from the evaluation and interviewer models to serve as a supervisory component.
Evaluation Topology. Prior benchmarks run each task independently and average the task metrics on separate subsets of the collected dataset to get an overall metric, neglecting the inherent dependencies between tasks while failing to support progressive and sustainable evaluation (Chang et al., 2024). In MACEval, we introduce four evaluation topologies, including line, star, tree, and hybrid networks, to establish the relationships among different evaluation streams. This also enables users to construct customized evaluation networks that better align with the assessment requirements of their target capability.
Figure 4: Performance curves and ACC-AUC values of different MLLM series. The red circles indicate the upper bound of the model’s capabilities. The left and right parts in subfigures depict the IQP and CU tasks, respectively.
Evaluating Continually. During the evaluation phase, an interviewer continually updates the data tuples ${({q,a})}_{t}$ based on new responses from the interviewee model, until it fails to answer correctly at a given level of difficulty. After that, the interviewee model, together with its performance records from the previous round, enters the next interview round. Such a spontaneous evaluation process goes beyond the limitations imposed by fixed dataset size, enabling deeper exploration of the model’s extreme performance while avoiding the manual collection of excessive and redundant data. Given the above objectives, we design an Area Under Curve (AUC)-inspired metric:
ACC-AUC $=\int_{a}^{\arg\mathop{\min}\limits_{t}\{t|{\text{ACC}}(t)=0\}}{{\text{ACC}}(t)dt,}$ (3)
${\text{ACC}}(t)$ $=\frac{\sum_{r=1}^{\mathcal{Q}_{total}}\mathbf{1}\left\{\mathcal{M}^{E}\left(q_{r,t}\right)\cong a_{r,t}\right\}}{\mathcal{Q}_{total}},$
where $a$ is the initial difficulty level, and $\mathbf{1}\left\{\cdot\right\}$ represents the indicator function, which returns 1 if the condition inside the braces holds, and 0 otherwise. $q_{r,t}$} and $a_{r,t}$} are the $r$-th query-answer pair at difficulty level $t$. $Q_{total}$ denotes the total number of query-answer pairs at a certain level. This simple and generally-applicable strategy enables us to evaluate the large models longitudinally and sustainably, as verified in our experimental analysis.
3.4 Utility-Capability Ratio
Building upon the single-dimensional evaluation strategy discussed previously, we introduce an evaluation routing-based metric to measure the overall utility-capability ratio of large models. Given an evaluation network with activated evaluation routing $\mathcal{P}$, the overall utility-capability ratio of the interviewee model $\mathcal{M}^{E}$ can be computed as follows:
$\epsilon_{\text{overall}}=\sum_{\left(\mathcal{M}^{E},\mathcal{M}^{I}\right)\in\mathcal{P}}\frac{\text{ACC-AUC}_{(\mathcal{M}^{E},\mathcal{M}^{I})}}{\arg\min_{t}\left\{t|\text{ACC}\left(t\right)=0\right\}}$ (4)
which quantifies the overall performance utilization under the limited capacity.
4 Experiments
4.1 Experimental Setup
Benchmark Candidates. Our experiments include 23 large models total, with a mix of cutting-edge proprietary models, open-source LLMs, and MLLMs. Specifically, for proprietary models, we include OpenAI models such as GPT-4o (OpenAI, 2024), and GPT-4.1 (2025-04-14) (Achiam et al., 2023), Google models such as Gemini-1.5-Pro, Gemini-2.0-Flash (Reid et al., 2024), and Gemini-2.5-Pro (Kavukcuoglu, 2025). For open-source LLMs, we include three mainstream language backbones widely used in many large models, i.e., the DeepSeeks (DeepSeek-V3 (Liu et al., 2024a) and DeepSeek-R1 (Guo et al., 2025)), the Qwens (Qwen3-8B (Team, 2025), Qwen2.5-{7, 14, 72}B (Team, 2024), Qwen2-7B (Yang et al., 2024a)), and the Llamas (Llama3.3-70B, Llama3.2-3B, Llama3.1-8B) (Grattafiori et al., 2024). For open-source MLLMs, we include models such as Qwen2.5-vl-{3, 7, 72}B (Bai et al., 2025), Qwen2-vl-7B (Wang et al., 2024), InternVL3-8B (Zhu et al., 2025), InternVL2.5-{8, 38}B (Chen et al., 2024b), InternVL2-8B (Chen et al., 2024c).
Figure 5: Utility-capability ratio comparison of five representative models across different dimensions.
Implementation Details. For all models, we perform evaluation using their respective templates under zero-shot settings. The initial difficulty level $a$ and the total number of queries in each level $\mathcal{Q}_{total}$ are set to 1 and 10, respectively. The max_new_tokens is set to the maximum value supported by each model to ensure the completeness of the response. We employ SSH connections to establish communication between the local interviewer model and the remote interviewee model running on a server. All experiments are conducted using a maximum of 8 RTX4090 24GB GPUs.
4.2 Main Results
In Tab. 1, Fig. 4, and Fig. 5, we analyze the performance of 15 LLMs and 8 MLLMs under different settings on 10 open-ended tasks. Our evaluation brings several important findings, as follows:
1) Identifying the gap in upper-bound performance via continual evaluation. As shown in Tab. 1, although the performance of different models varies, models of the same scale or generation generally share a similar upper limit of performance. This phenomenon is particularly evident in arithmetic and algorithmic tasks that require strong reasoning abilities, reflecting the common performance bottleneck of current LLMs.
2) Models with downloadable weights currently lag behind the top-performing models. All five proprietary models, especially the more advanced Gemini 2.5 Pro and GPT-4o, have shown superior performance to open-source models, such as the Qwen 2.5 and Llama 3.3 series, on most tasks. It is worth noting that DeepSeek-R1, with its powerful reasoning capabilities, outperforms proprietary models on over 60% of the tasks.
Figure 6: Pearson correlation between different tasks.
Figure 7: Maximum performance level vs. ACC-AUC. We plot the performance of 15 LLMs compared to the corresponding maximum difficulty levels, with each a fit line and correlation coefficient. The red dotted lines denote the performance saturation line.
3) Scaling law vs. Technological advancement. Tab. 1 and Fig. 4 show a notable performance improvement in the evaluated capabilities when comparing larger models to smaller ones, as well as newer model generations to their predecessors, showing a relatively strong scaling law. Meanwhile, the latest Qwen3-8B benefited from its thinking mode, significantly surpassing the previous Qwen2.5-14B, and almost on par with Qwen2.5-72B in math and coding.
4) Performance utilization varies across models and tasks. As shown in Tab. 1, DeepSeek-R1 has the highest overall utility-capability ratio. While the Qwen2.5 series surpasses its next generation Qwen3-8B, indicating the possible instability resulting from the combination of the reasoning mode. Fig. 5 exhibits the discrepancies in performance utilization among tasks, where most models show low correlation between their performance utilization rate and absolute performance ranks, suggesting another optimization direction for large models.
5) Steep decline in low-level visual perception. In Fig. 4, we observe a sharp performance cliff in the low-level IQP tasks, which suggests potential Just Noticeable Difference (JND) points in the vision backbones of MLLMs. Besides, the InternVL series exhibits more sensitive perceptual responses compared to the Qwens and other models.
4.3 Correlation Analysis
In Fig. 6, we compute the Pearson correlation coefficient among all pairs of tasks. Results show that, unsurprisingly, text comprehension, algorithm reasoning, and coding all correlate with one another (avg. $r>0.81$). Moreover, math tasks involving intensive numerical computation correlate relatively weakly with all other categories, which, together with the extremely low maximum capability level reported in Tab. 1, indicates the uniqueness of this capability. We further investigate the relationship between maximum capability level and long-term ACC-AUC in Fig. 7, where two dimensions are highly correlated (avg. $r\approx 0.888$) in all tasks, and few models reach saturated performance within a limited capability level.
4.4 Effect of Different Evaluation Topologies
1) Ablation Study on Interviewer Diversity. Here, we discuss the performance differences under different interviewer settings. As listed in Tab. 2, the evaluation results remain consistent across different interviewers in most scenarios. However, when evaluating tasks that involve more complex text generation, using a weaker model as the interviewer may limit the exploration of the interviewee model’s upper-bound capabilities.
Table 2: Ablation study on different interviewers.
Interviewee Interviewer STUmask. Arith.scale BTT Debug.
Gemini 1.5 Pro Gemini 2.5 Pro 5.218/6 0.8/1 2.8/3 14.9/17
DeepSeek-V3 5.087/6 1.4/2 3.3/4 14.1/15
Qwen2.5-7B 5.083/6 1.3/2 2.6/3 12.9/13
Qwen3-8B Gemini 2.5 Pro 3.095/4 1.0/2 1.4/2 11.2/14
DeepSeek-V3 2.818/4 0.9/2 1.4/2 11.7/14
Qwen2.5-7B 2.905/4 1.0/2 2.0/3 10.9/12
Llama3.1-8B Gemini 2.5 Pro 2.946/4 0.1/1 0/0 9.6/12
DeepSeek-V3 2.930/4 0.3/1 0/0 9.4/12
Qwen2.5-7B 2.998/4 0.2/1 0/0 9.0/11
Table 3: Results on different role settings. Triple means Qwen2.5-7B, GPT-4o, and Claude 3.7 Sonnet are employed as interviewers.
Interviewee Interviewer Supervisor STUdisrupt. Arith.scale BTT
Qwen2.5-7B GPT-4o Claude 3.7 Sonnet 6.797/8 0.9/1 1.6/2
Qwen2.5-7B Qwen2.5-7B 6.833/8 1.0/2 1.6/2
Triple Qwen2.5-7B 6.802/8 0.9/1 1.6/2
GPT-4o GPT-4o Claude 3.7 Sonnet 8.855/10 1.7/2 3.1/4
GPT-4o GPT-4o 8.855/10 1.7/2 3.1/4
Table 4: Error rate of w/ and w/o message passing (GPT-4o as interviewer). End nodes are in bold.
Topology STUdisrupt. STUmask. SP Arith.scale Arith.oper. BTT SPSnode SPSedge Generation Debugging
1-hop line (w/o) 36% 42% 18% 2% 50% 52% 26% 36% 30% 18%
[…]
2-hop line (w/) 18% (-24%) 4% (-46%) 22% (-14%) 14% (-4%)
2) Multiple Roles Exploration. We further evaluate the effect of different role settings, where the same model is used simultaneously as the interviewee, interviewer, and supervisor. Concretely, we test Qwen2.5-7B and GPT-4o in three tasks, i.e., STUdisrupt., Arith.scale, and BTT. In Tab. 3, we observe that changing the supervisor does not explicitly affect the resulting performance. Meanwhile, after changing the interviewer from GPT-4o to Qwen2.5-7B, the evaluated performance of Qwen2.5-7B on STUdisrupt. and Arith.scale tasks grows slightly. We suspect that this is due to the limited diversity of the generated queries stemming from the relatively weak capability of the interviewer model (similar observation as Tab. 2). Therefore, we employ a 1v3 star evaluation topology to solve the mild variance for the same interviewee model across interviewers, which indeed improves the reliability of the evaluation process. Moreover, we find a high error rate (56%) in the BTT task, calculated by the misalignment between the external function and self-generated ground-truth, when Qwen2.5-7B serves as the interviewer. This result is much lower in GPT-4o’s setting, showing the necessity of using relatively powerful models and incorporating external functions.
3) Overall performance comparison based on the evaluation routing. In Fig. 13, we show the sum of normalized performance for 15 LLMs based on the given evaluation routes. We notice that proprietary models still maintain a dominant position. The characteristic of $\epsilon_{\text{overall}}$ lies in considering underlying inter-task dependencies, incorporating information-aware evaluation routing instead of independently assessing and aggregating task scores. We then compare these results to other mainstream benchmarks (LiveBench (White et al., 2025) and ChatBot Arena (Chiang et al., 2024)), shown in Fig. 9. It can be observed that the results of MACEval highly correlate with mainstream benchmarks (avg. $r=0.9509$), demonstrating the effectiveness of MACEval and the scalable design concept of MACEval, where models can conduct fair performance tests within a fixed evaluation network.
Figure 8: Error rate (%) versus propagation round.
4.5 Potential Error Propagating
1) Message passing improves the quality of question generation. We focus on two types of question generation errors caused by hallucinations from the interviewer model during the evaluation process, i.e., format violations and redundant questions. Tab. 4 demonstrates the effectiveness of the message passing mechanism in improving the question quality between two related tasks with an average error rate reduction of 22%. For MLLM evaluations, since the subject of the visual perception tasks is the image itself, no obvious mistakes in question generation were detected.
2) Impact of Proactive Error Injection. We conduct extra experiments to explore the robustness of MACEval framework. Specifically, we add interference characters (unordered and meaningless strings) into the message, which are then taken to the next node. We evaluate the data generation process of the binary tree traversal (BTT) task, which has the highest initial error (52%, Tab. 4). The results in Fig. 8 show that injecting irrelevant textual contents exhibits little influence on the inherent error rate that depends on the difficulty of the task itself. Conversely, the progressive accumulation of erroneous content renders the information more readily detectable by a third-party supervisor, facilitating its timely suppression and the issuance of a regeneration directive to the interviewer model.
Table 5: Results on different numbers of queries. ACC-AUC/Max. Perf. is reported.
Task\#Query 10 20 30 50
STUmask. 4.247/5 4.256/5 4.253/5 4.244/5
BTT 2.40/3 2.50/3 2.43/3 2.46/3
Table 6: Results of the statistical test.
#Query@N times STUmask. BTT
10@10 [4.2444, 4.2558] [2.4122, 2.4618]
20@10 [4.2457, 4.2533] [2.4265, 2.4570]
pp-value 0.8486 0.7227
Figure 9: The performance of LLMs on different benchmarks, compared to a best-fit line. We compare the evaluation results in relative performance on MACEval vs. LiveBench, and MACEval vs. ChatBot Arena.
4.6 Stability Analysis
To validate the stability of evaluation data generation, we compare the performance differences in terms of the number of generated queries at the same difficulty level on STUmask. and BTT tasks, as shown in Tab. 5. The interviewee and interviewer are Qwen2.5-72B and GPT-4o, respectively. We also conduct two-sided t-test and reported 95% PI and $p$-value for 300 queries in Tab. 6. We can observe that there is no significant difference between the results of 10@10 and 20@10, as the $p$-value>>0.05. Their 95% PIs have a considerable overlap, indicating the robustness of the evaluation process. Furthermore, the 95% PI of 20@10 is more concentrated than that of 10@10, showing the precision of increasing the number of rounds under the given difficulty level. See more experimental results in the Appendix. C.
4.7 Comparison to Other Large Model Benchmarks
We compare the normalized evaluation results obtained by MACEval to popular benchmarks (LiveBench (White et al., 2025) and ChatBot Arena (Chiang et al., 2024)), shown in Fig. 9. It can be observed that the results of MACEval highly correlate with mainstream benchmarks (avg. $r=0.9509$), demonstrating the effectiveness of MACEval and proving the potential of in-process problem generation.
5 Conclusion
This work introduces MACEval, a novel dynamic evaluation framework for large generative models, which provides effective solutions to avoid data contamination and reduce human participation by a multi-agent system-based automatic evaluation network. To identify the upper bound of performance and evaluate continually, we propose an AUC-inspired metric, further incorporating the evaluation topologies to measure the overall performance of large models. Experiments on five high-profile capabilities demonstrate the effectiveness, efficiency, and flexibility of MACEval compared to existing benchmarks, presenting a promising direction for achieving fair evaluation.
Impact Statement
Our work holds the potential for significant societal impact, both positive and negative. First, the evaluation is normally considered as important as training in developing large models. An insightful evaluation can offer promising directions to large model developers. The main motivation of this work is to build an automatic, scalable, and trustworthy evaluation framework for large models, aligning closely with the rapid iteration of current models. Our MACEval offers a promising direction that using large models to tackle the evaluation problems for themselves has lower costs than the traditional human-dependent form on broad occasions. Second, the proposed interview-like autonomous evaluation pipeline with open-ended settings is more conducive to exploring the performance boundaries of large models with less subjective bias. Third, although our research is not directly related to the design of generative models, the possibility of spontaneously generating NSFW content in other tasks due to hallucinations or ambiguity in the content of the dialogue cannot be ruled out, since we adopted a human-free autonomous multi-agent collaboration framework in the evaluation process, where the evaluation content is autonomously generated.
Appendix A Detailed Open-Ended Tasks Construction
The objective of curating open-ended tasks for large model evaluation is to mitigate data contamination issues. We design 9 tasks based on fully AI-generated and difficulty-controllable principles, spanning five capabilities, including visual perception, text comprehension, math, algorithm reasoning, and coding, to preliminarily validate the effectiveness of our MACEval framework:
  • Image Quality Perception (IQP): We select a set of keywords (people, animals, nature, plants, cars, objects, buildings, textures, sports, and interiors) based on the categorization of the renowned image website Unsplash [^1] https://unsplash.com/, which served as the source for generating image prompts. Then, we adopt an interviewer-assisted prompt generation strategy to obtain $\mathcal{Q}_{total}$ prompts. Five prevailing text-to-image models, including Stable Diffusion-1.5 (Rombach et al., 2022), Stable-Diffusion-3.5-medium (AI, 2024), Playground v2.5 (Li et al., 2024b), FLUX.1-schnell (Labs, 2024), and Sana (Xie et al., 2024), are randomly invoked externally to generate original image contents. The output image size is set to $512^{2}$, and the num_inference_steps is set to the default value. As for the difficulty control, we follow the adjustable parameters and ranges in Image Adjustment Tool Palette (Canon, 2024) and generate a set of Gaussian white noise by setting the variance to $1:\infty:1]$ (which represents values starting from $1$ with a step size of $1$) and add it to the original images.
  • Content Understanding (CU): This task evaluates the model’s ability to distinguish image content. We download over 100 different source icons from FLATICON [^2] https://www.flaticon.com/ and randomly form part of them into a grid image, which explicitly avoids image content contamination, as shown in Fig. 10. Then, we examine whether the interviewee model can correctly answer the exact number of target icons in the given grid images. The task difficulty is controlled by the number of grids, i.e., the more grids there are, the more interference terms are presented. We begin with the $3\times 3$ size.
  • Scrambled Text Understanding (STU): Humans can grasp the general meaning of sentences and paragraphs even with some misspelled words, without significantly affecting overall comprehension. However, whether LLMs can achieve the same remains an open question. To create the questions for this task, we instruct the interviewer model to generate a text of approximately 100 words (e.g., story, diary, or prose), which is then processed by external functions to apply two types of perturbations: 1) disrupt the character order within words. For each text, we flip a certain proportion ($N^{word}_{disp}/N^{word}_{total}$) of correctly spelled words to misspelled words (White et al., 2025), where $N^{word}_{disp}$ and $N^{word}_{total}$ denote the number of disrupted words and the total number of words, respectively; 2) mask a certain proportion ($N^{char}_{masked}/N^{char}_{total}$) of characters with ‘*’, where $N^{char}_{masked}$ and $N^{char}_{total}$ denote the number of masked characters and the total number of characters, respectively. The resulting text is then sent to the interviewee model for completion and comprehension.
  • String Parsing (SP): To investigate the discernibility of LLMs in processing long texts, we design a needle-in-a-haystack evaluation task, requiring the LLMs to count the occurrences of a target character within a long string. We perform similar instructions to the STU task for string generation and add the needle ‘-’ to a random position within the string. Then, the string length is incrementally increased based on the interviewee’s performance to elevate task difficulty.
  • Arithmetic: Considering the inherent challenge of quantifying the difficulty of mathematical problems, which is normally associated with factors such as problem complexity, level of abstraction, the number of solution steps, and the methods employed to solve it, we select two basic arithmetic problems with quantifiable difficulties: 1) multiplication of decimals with increasing digits, and 2) operations with increasing number of terms.
    Figure 10: Visualization of grid images with different icons in the content understanding task.
  • Binary Tree Traversal (BTT): The binary tree is a common data structure widely used in computer science and programming. In general, a binary tree structure can be uniquely determined by given its pre-order and in-order traversals. Based on this, we prompt the interviewer model to generate a binary tree with its pre-order and in-order traversals, then query the interviewee model to output the post-order traversal. The number of nodes is used to control the task difficulty, namely, the structural complexity of a binary tree.
  • Shortest Path Search (SPS): The shortest path algorithm is an important concept in graph theory used to find the shortest path between two nodes in a weighted graph. Taking the famous Dijkstra algorithm as an example, its time complexity is $O(V^{2})$ and $O((V+E)\log V)$ when using a regular array and a priority queue, respectively. Inspired by this, we first instruct the interviewer model to generate a graph with an adjacency matrix along with the start and end nodes, which is further sent to the interviewee model for shortest path computation. The number of nodes and edge density are used to control the difficulty.
  • Code Generation: Similar to the mathematical problems, the difficulty of coding problems is also hard to define. In this paper, we focus on the cyclomatic complexity, which is a software metric that quantifies the complexity of a code’s control flow, specifically measuring the number of linearly independent paths through the code. We first instruct the interviewer model to generate an initial code demand and then send this to the interviewee model for code generation. After that, the generated code will be sent back to the interviewer model for examination. If the interviewee completes the current task, a new requirement will be randomly added, such as incorporating a ‘for’ loop, a ‘while’ loop, or an ‘if’ conditional statement. This process will continue iteratively until the interviewee fails to produce the correct output.
  • Code Debugging: As for the debugging task, we adopt the same code generation demand to the interviewer model and then mask or delete a certain proportion of them by an external function. The interviewee model is asked to repair the missing sections.
Note that for arithmetic, BTT, and SPS tasks, we implement external code-based supervision to prevent potential generation errors made by the interviewers themselves, shown in Fig. 12.
Figure 11: Example of four typical evaluation topologies as well as their variants.
Figure 12: External upervision functions used in the BTT and SPS tasks.
Appendix B Additional Information of Evaluation Topology
We distinguish evaluation topologies into four primary categories: line, star, tree, and hybrid network, as shown in Fig. 11.
  • Line: Line-based evaluation topology is the simplest structure where nodes (such as interviewee models and interviewer models) are connected sequentially along a single communication line. In other words, in each round of task evaluation, only one interviewer model engages in dialogue-based assessment with the interviewee model.
  • Star: Star topology means that one interviewee model is connected to multiple interviewer models, with a maximum hop count of 1. The typical assessment scenarios corresponding to this evaluation topology include: a) $N$ interviewers jointly assess the performance under a specific task and conduct a comprehensive performance discussion or merely average the score; b) for each interviewee, all tasks or capabilities are evaluated in parallel. Each line within a star topology denotes a single evaluation process.
  • Tree: Tree-based evaluation topology possesses an innate hierarchical structure that enables decomposable evaluations from low-level abilities to high-level abilities. For example, when evaluating the math capability, one can start with the fundamental arithmetic tasks and proceed to more complex calculus, equation solving, and mathematical proof tasks.
  • Hybrid Network: Compared to other single-type topologies, the hybrid network presents a more intricate web of node relationships, where a new edge type (interviewee-interviewee) and a cyclic structure are introduced. Specifically, such a structure is more suited to a panel discussion form of evaluation task. For example, evaluating the collaborative communication skills and debating abilities of large models for more truthful answers, when confronted with the task of designing solutions to complex problems (Khan et al., 2024).
Figure 13: Overall performance comparison among 15 LLMs based on the right evaluation network, where colored clusters represent evaluation agents for different capabilities and red arrows denote activated evaluation routes.
Appendix C Additional Experimental Results
Subjective Alignment of the Supervisor. We conducted extra subjective experiments where we quantified their alignment with human annotations. Specifically, we adopt two schemes: ‘yes-or-no’ evaluation and scoring evaluation (Chen et al., 2024a). 2 experts (with experience in publishing articles on large models) and 8 public participants from the campus are invited to participate. We selected four tasks with high error rates (according to Tab. 4), including STUdisrupt., STUmask., Arith.oper., and BTT, for evaluation. We use Claude 3.7 Sonnet and Gemini 2.5 Pro as the supervisor. GPT-4o is the interviewer. Note that in these tasks, we have already implemented external algorithmic function checks for the ground-truth errors (e.g., “23445 + 22784 = ?”). Therefore, human and third-party supervisory models primarily focus on errors related to textual hallucinations. In ‘yes-or-no’ evaluation, the human participants and the supervisor model are instructed to answer “whether there exists any hallucination in the evaluation dialogues? Return yes or no.” In the scoring evaluation, we instruct “Please carefully judge the quality of the evaluation process dialogue based on the task requirements,” and provide a detailed scoring standard with a 1-5 Likert scale:
  • Level-1: The response is accurate and trustworthy
  • Level-2: The response is mostly correct but contains minor, non-core factual errors
  • Level-3: A mix of fact and fabrication makes the response partially unreliable
  • Level-4: The core claim is false, or the response is mostly fabricated and misleading
  • Level-5: The response is almost entirely invented, nonsensical, or detached from reality.
In Tab. 7, we observe a significantly high correlation between human and the supervisor models in both settings, demonstrating their effectiveness in supervising the evaluation process. In addition, we found that over 95% of the queries received a rating of 4 or higher, showing the effectiveness of the joint generation scheme of the interviewer model and the external function.
Table 7: The Spearman’s $r$ between the score of the model and all raters for ‘scoring’ setting. The human agreement percentage in the form of experts/crowds for a ‘yes-or-no’ setting.
Setting Supervisor STUdisrupt. STUmask. Arith.oper. BTT
Scoring Claude 3.7 Sonnet 0.8419 0.8593 0.8862 0.8122
Scoring Gemini 2.5 Pro 0.8522 0.8668 0.8912 0.8217
Yes-or-No Claude 3.7 Sonnet 0.944/0.923 0.935/0.920 0.936/0.918 0.903/0.898
Yes-or-No Gemini 2.5 Pro 0.948/0.929 0.938/0.922 0.941/0.927 0.925/0.908
Compare ACC-AUC to Static Evaluation Baselines with Difficulty Settings. We conduct additional experiments for comparing performance trends across easy, medium, and hard tasks in static benchmarks under the framework of MACEval. Specifically, we use the problems from E2H-AMC, a subset in Easy2Hard-Bench (Ding et al., 2024), where the difficulty rating is represented by the percentage of students who answered each question correctly. Here are three problem examples:
  • Cagney can frost a cup cake every <20><20> seconds and Lacey can frost a cupcake every <30><30> seconds. Working together, how many cupcakes can they frost in <5><5> minutes?—Avg. Difficulty = 0.134
  • Find the number of pairs of integers (a, b) with <1><1> $\leqslant a<b\leqslant$ <57><57> such that $a^{2}$ has a smaller remainder than $b^{2}$ when divided by <57><57>.—Avg. Difficulty = 0.587
  • In a $16>\times<16>$ table of integers, each row and column contains at most <4><4> distinct integers. What is the maximum number of distinct integers that can be in the whole table?—Avg. Difficulty = 0.784
We uniformly sample 10 question templates with incremental difficulty rating from 0.1035 to 0.8548 (The larger the value, the higher the difficulty). During the evaluation data generation, the interviewer model with an external function-assisted is instructed to generate a query-answer by changing the values in ‘<><>’. We set the number of queries at each difficulty level to 100. Rather than relying on manually annotated fine-grained difficulty scores, we increase the number of testing rounds within each predefined difficulty level to obtain a more precise performance estimate. Here, we tested two models: Qwen 2.5-VL-72B and Gemini 1.5 Pro. GPT-4o is used as the interviewer model. In Tab. 8, we observe that the overall declining trend on both dynamic and static settings is similar. The performance under MACEval with multi-round queries exhibits slightly higher values than those with static settings, showing a more genuine performance while demonstrating its potential superiority for longitudinal evaluation against data contamination.
Table 8: Data points in discrete form. The upper and bottom parts are the accuracy results under MACEval and static settings, respectively.
Difficulty level 1 2 3 4 5 6 7 8 9 10
Qwen2.5-72B 0.57 0.42 0.32 0.27 0.24 0.18 0.14 0.10 0.10 0.09
Gemini 1.5 Pro 0.51 0.36 0.24 0.21 0.16 0.14 0.10 0.08 0.08 0.08
Qwen2.5-72B 0.53 0.43 0.29 0.21 0.18 0.17 0.12 0.10 0.08 0.09
Gemini 1.5 Pro 0.46 0.39 0.22 0.12 0.10 0.07 0.06 0.04 0.04 0.03
Table 9: Results on the expanded language understanding tasks.
Model Interviewer Task1-1 Task1-2 Task2
GPT-4o GPT-4o 54.4/88 41.3/45 57.5/93
Gemini 2.0 Flash 67.2/113 40.8/44 70.3/116
Qwen2.5-72B 49.8/75 41.9/45 51.6/77
Qwen2.5-7B 21.5/26 29.3/38 21.8/27
GPT-4o Claude 3.7 Sonnet 53.8/87 41.9/45 58.9/95
Gemini 2.0 Flash 65.8/109 42.7/46 74.4/119
Qwen2.5-72B 50.3/77 41.7/45 53.4/78
Qwen2.5-7B 21.9/26 29.5/39 22.4/28
GPT-4o GPT-4o, Claude 3.7 Sonnet, Gemini 2.5 Pro, DeepSeek V3 55.6/88 41.7/45 58.2/95
Gemini 2.0 Flash 67.1/112 41.7/45 73.8/117
Qwen2.5-72B 50.8/76 41.7/45 52.9/77
Qwen2.5-7B 21.8/26 29.6/39 21.9/27
Table 10: Results on the equation solving task.
Model Equation Solving
GPT-4o 1.4/2
Gemini 2.5 Pro 2.1/3
Gemini 2.0 Flash 1.4/2
Gemini 1.5 Pro 1.1/2
Qwen2.5-7B 0.8/1
Qwen2.5-14B 0.8/1
Qwen2.5-72B 1.3/2
Llama3.1-8B 0.6/1
Results on Tasks that are Less Structured and Quantifiable. We further expanded the task suite to include language understanding tasks that are less structured or quantifiable. Specifically, we focus on the following tasks. 1) External material: To avoid data contamination, we manually collected 10 recent news articles from world economic forum [^3] https://www.weforum.org/ (time period: 2025.7.20-7.25, later than the cutoff date of the evaluated models) to ensure that they are not in the training data. 2) Self-generated: The interviewee model is required to generate text contents with a certain length and requirements. Below we provide detailed settings.
  • Task 1-1 (External): The interviewer randomly requires the interviewee model to paraphrase, simplify, or summarize the given text content with a certain output format. We score tasks purely by their adherence to the instructions (White et al., 2025). The interviewer models are instructed to increase the number of queries in each round until the finishing rate equals a very low value. However, in these experiments, we found a relatively slow rate of decline. It is time-consuming to reach the scenario with 0% finishing rate compared to other tasks. Therefore, we report the ACC-AUC at 50% finishing rate and the corresponding number of queries, which can also validate the feasibility of MACEval in handling open language understanding tasks.
  • Task 1-2 (External): We further employ the language bootstrapping (Yang et al., 2024c) (e.g., the proportion of irrelevant words or sentences) to adjust the difficulty rate each success round while satisfying the principle of MACEval (pushing to the limit). Therefore, we report the ACC-AUC and the bootstrapping proportion (%).
  • Task 2 (Self-generated): The interviewer models are instructed to increase the number of queries (difficulty rate) in each round until the finishing rate equals a very low value. An external Python function is deployed to count word length for accuracy calculation. Same with scenario 1-1, we report the ACC-AUC at 50% finishing rate and the corresponding number of queries due to the slow decline rate.
As shown in Tab. 9, four models are selected for evaluation, and three different evaluation topologies are investigated (including 1-hop line and 1v4 star). We find a tiny difference in ACC-AUC/Max. Perf. when changing GPT-4o to Claude 3.7 Sonnet under 1-hop line. This can be attributed to the small disparity of the in-process data, which can be mitigated under the 1v4 star evaluation topology, where a mean opinion score (MOS)-based strategy is adopted, showing good robustness and flexibility of the MACEval framework. We notice the results in Task 1-1 and Task-2 are similar, indicating that the influence of the data source in certain tasks is negligible and validating the suitability of AI-generated text in such tasks.
More Evaluation on Math Tasks. We expand the math tasks to solving the equations, and choose 8 models for evaluation, where GPT-4o serves as the interviewer model. The difficulty level is set to the number of variables. We require it to retain four decimal places. As listed in Tab. 10, we observe that those open-source models, even with large-scale parameters or mainstream proprietary models, such as GPT-4o and Gemini1.5 Pro, can only solve equations with two variables, showing great potential for general large models to improve. Surprisingly, the cutting-edge Gemini 2.5 Pro, a model skilled in mathematics and reasoning, is only capable of solving systems of three linear equations in three variables to a limited extent.
Figure 14: The working mechanism of the third-party supervisor involved along the evaluation pathway and responsible for validating the entire evaluation network.
C.1 Visualization of the Evaluation Stream
To illustrate the working mechanism of MACEval more intuitively, we visualize the dialogue example of the third-party supervisor and prompt templates for both interviewee and interviewer in Fig. 14 and Appendix G. In this paper, we adopt the Claude 3.7 Sonnet (Anthropic, 2024) as the supervisor.
Appendix D Limitations
There are two potential limitations acknowledged below. First, we apply evaluation topologies with fewer hops, i.e., nodes and edges, to preliminarily demonstrate the effectiveness of MACEval. In the future, it is possible to construct larger evaluation networks to include a richer array of capability assessments. Second, in MACEval, the evaluation data stream is composed of question-answer pairs graded by difficulty. One of the principles in constructing such open-ended tasks is that the difficulty should be quantifiable, which imposes limitations on the implementation of many tasks.
Appendix E Future Possibilities
Apart from the comments on possible next steps for research related to the evaluation of large models that have already been provided, this section is devoted to extending some of them and exploring additional topics with good potential based on our understanding and rethinking of the field.
  • Focus on Efficient Evaluation. Currently, most evaluations are trending towards larger scales and broader scopes, which significantly exacerbate the costs associated with data collection and maintenance. Therefore, the development of efficient evaluation pathways is of great importance. For instance, it may be possible to assess a large model’s capabilities in a particular dimension by evaluating a small number of key samples. Conducting principal component analysis and dimensionality reduction on evaluation data to identify efficient and rapid evaluation routes could emerge as a significant research direction.
  • Evaluating the evaluations. Evaluations are critical for understanding the capabilities of large models. Its fairness, comprehensiveness, reproducibility, timeliness, scalability, transparency, and practicality are key points to serve as a ”good” evaluation. Recently, with the rapidly increasing number of various benchmarks, there is a need for objective metrics to evaluate evaluations themselves.
It is hoped that these can provide actionable ideas for researchers and would trigger further discussion, and more importantly, new exploration in this area.
Appendix F Usage of Large Language Models
We acknowledge that Large Language Models (LLMs) were used in the preparation of this manuscript solely for grammatical corrections and sentence polishing
Appendix G Prompt Templates
G.1 Prompt Templates for Visual Perception Tasks
Example for Interviewer:
<|System Prompt|>
You are a helpful assistant for prompt generation in text-to-image tasks.
<|User Prompt|>
Please generate a prompt regarding the generation of {keyword} images. Return text in <text></text> format.
<|Interviewer’s Answer|>
<text>A high-resolution natural landscape at sunrise, viewed from a hilltop. The foreground shows dewy grass and wildflowers gently swaying in the breeze. In the middle ground, a winding dirt path leads through a lush green forest with scattered rays of sunlight filtering through the trees. In the background, soft morning mist blankets a valley with a river glistening under the golden light of the rising sun. Birds are flying low near the horizon, and the sky transitions from deep orange to pale blue.</text>
<|External image generation model|>
def SD3_5 (prompt):
<|External distortion algorithm|>
def add_noise (img, noise_level):
<|Resulting content|>
Reference image: <image1>
Distortion image: <image2>
Noise level: <noise_level>
Example for Interviewee:
<|User Prompt|>
Given two images <image1> and <image2>, please answer if there are any differences in noise between them. Return judgment <yes> or <no> with analysis in format:
<Judgment>xxx</Judgment>
<Analysis>xxx</Analysis>
<|Interviewee’s Answer|>
<Judgment>No</Judgment>
<Analysis>There are no differences between these two images. They appear to be identical.</Analysis>
G.2 Prompt Templates for Content Understanding Tasks
Example for Interviewer:
<|System Prompt|>
You are a helpful assistant.
<|User Prompt|>
Please generate two random numbers from 1 to 100 and from 1 to {{level+2} $\times$ {level+2}}, respectively. Return answer in the following format:
Number1: <> # The first one denotes the selected icon’s ID.
Number2: <> # The second one denotes the number of the same icons.
<|Interviewer’s Answer|>
Number1: <67>
Number2: <11>
<|External grid_image generation algorithm|>
def grid_image (image_id, size):
<|Resulting content|>
Icon image: <image1>
Grid image: <image2>
Grid size: <level+2>
Example for Interviewee:
<|User Prompt|>
Given an icon image <image1> and a large grid image <image2>, please answer how many icons <image1> are there in the grid image <image2>? Return number only in format:
Number: <>
<|Interviewee’s Answer|>
Number: <4>
G.3 Prompt Templates for Scrambled Text Understanding Tasks
Example for Interviewer (Disruption):
<|System Prompt|>
You are an expert in the field of literature, specializing in writing stories, journals, and essays.
<|User Prompt|>
Please generate a piece of text, about 100 words, which can be a story, journal, or essay. Output the text between the <text></text> symbols.
<|Interviewer’s Answer|> <text>
Amid the dense forest, Lena stumbled upon an ancient, moss-covered cabin. The air was thick with the scent of pine and mystery. Inside, a dusty journal lay open on a wooden table, its pages filled with cryptic symbols and sketches of strange creatures. As she flipped through the pages, a chilling wind howled through the cracks, and the door creaked open. A shadowy figure loomed in the doorway, eyes gleaming with eerie familiarity. Lena’s heart raced as the figure whispered, “You’ve finally found it.” The journal’s last page read, “The key to the past is in the present.”
</text>
<|Adding Perturbations (Disrupt the character order)|> def disrupt_words (ori_text, disp_ratio):
<|Resulting Text|>
Amid the dnese foerst, Lena stubmled upon an nacient, moss-covered cabin. The air was tihck with the scent of pnie and msytery. Inside, a dusty jorunal lay open on a wooedn table, its paegs filled with cryptic syombls and skecthes of starnge crateures. As she flipped thurogh the pages, a chilling wnid howled through the cracks, and the door crekaed open. A sahdowy fiugre loomed in the doroway, eyes gleaming with eerie famiilarity. Lena’s hraet raced as the fiugre whispered, “You’ve fainlly found it.” The journal’s last page read, “The key to the past is in the present.”
Example for Interviewee (Disruption):
<|User Prompt|>
Here is an text with typos: <text>...</text>. Please output this exact text, with no changes at all except for fixing the misspellings. Please leave all other stylistic decisions like commas and US vs British spellings as in the original text. Output the revised text between the <text></text> symbols.
<|Interviewee’s Answer|> <text>
  Amid the dense forest, Lena stumbled upon an ancient, moss-covered cabin. The air was thick with the scent of pine and mystery. Inside, a dusty journal lay open on a wooden table, its pages filled with cryptic symbols and sketches of strange creatures. As she flipped through the pages, a chilling wind howled through the cracks, and the door creaked open. A shadowy figure loomed in the doorway, eyes gleaming with eerie familiarity. Lena’s heart raced as the figure whispered, “You’ve finally found it.” The journal’s last page read, “The key to the past is in the present.”
</text>
Example for Interviewer (Mask):
<|System Prompt|>
same as the above
<|User Prompt|>
same as the above
<|Interviewer’s Answer|>
same as the above
<|Adding Perturbations (Mask characters with * )|>
def mask_characters (ori_text, mask_ratio):
<|Resulting Text|>
Amid the de*se fo**st, Lena stu**led up*n an an**ent, moss-cov*red cabin. The a*r was thi*k with the scent of p*ne and mys*ery. In*ide, a *usty jour*al lay op*n on a **oden ta*le, its pa*es filled with cryptic s**bols and sket**es of st**nge cre*tu*es. As she fl**ped through the p*ges, a chilling w**d howled thro*gh the c*acks, and the door creaked open. A sh***wy f*gure loomed in the d**rway, eyes gleaming with eerie famil**rity. Lena’s h*art raced as the fi**re whis*ered, “You’ve fina**y found it.” The jou**al’s last page read, “The key to the past is in the pr*sent.”
Example for Interviewee (Mask):
<|User Prompt|>
Here is an text with masks (*): <text>...</text>. Please output this exact text, with no changes at all except for filling the masks. Please leave all other stylistic decisions like commas and US vs British spellings as in the original text. Output the revised text between the <text></text> symbols.
<|Interviewee’s Answer|> <text>
  Amid the dense forest, Lena stumbled upon an ancient, moss-covered cabin. The air was thick with the scent of pine and mystery. Inside, a dusty journal lay open on a wooden table, its pages filled with cryptic symbols and sketches of strange creatures. As she flipped through the pages, a chilling wind howled through the cracks, and the door creaked open. A shadowy figure loomed in the doorway, eyes gleaming with eerie familiarity. Lena’s heart raced as the figure whispered, “You’ve finally found it.” The journal’s last page read, “The key to the past is in the present.”
</text>
G.4 Prompt Templates for String Parsing Tasks
Example for Interviewer:
<|System Prompt|>
You are a helpful assistant proficient in generating text.
<|User Prompt|>
Please generate a {level}-character-long string that may include English letters (both uppercase and lowercase) and special characters such as: !!, @@, #, %, &. Then, insert four ‘-’ characters at random positions in the string. Return string in String: <> format.
<|Interviewer’s Answer|>
String: <a-D#fG%kL-qW!!zXe@@R-tY&>
<|External algorithm supervision|>
def task_calibration (ori_text):
<|Resulting Text|>
String:<a-D#fG%kL-q-W!!zXe@@R-tY&>
Example for Interviewee:
<|User Prompt|>
Here is a long string: <text>...</text>. How many ‘-’ characters are there in this string? Return the number in the following format:
Number: < >
<|Interviewer’s Answer|>
Number: <4>
G.5 Prompt Templates for Arithmetic Tasks
Example for Interviewer (Scale):
<|System Prompt|>
You are a helpful assistant proficient in mathematics.
<|User Prompt|>
Please generate a multiplication problem with the correct answer involving two n-digit decimal numbers. The two numbers must be different from those used in previous problems.
Return question and correct answer in the following format:
<question> </question>
<answer> <answer>
<|Interviewer’s Answer|>
<question> What is 123.456 × 789.123? </question>
<answer> 97406.100088 </answer>
<|External algorithm supervision|>
def task_calibration (ori_text):
<|Resulting Text|>
<question> What is 123.456 × 789.123? </question>
<answer> 97421.969088 </answer>
Example for Interviewee (Scale):
<|User Prompt|>
Please solve the given math problem and return your answer in <answer> </answer> format.
<|Interviewer’s Answer|>
<answer> 97461.969 </answer>
Example for Interviewer (Operation):
<|System Prompt|>
You are a helpful assistant proficient in mathematics.
<|User Prompt|>
Please generate an arithmetic problem with the following requirements:
1. Each number should have exactly {digit} significant digits
2. The numbers should be randomly generated, not following any pattern
3. The problem should use exactly {level+1} operators such as (+,,×,/)(+,-$\times,$/)
4. You can use parentheses () freely to group operations
5. The numbers should be different from the former problems
Return question and correct answer in the following format:
<question> </question>
<answer> <answer>
<|Interviewer’s Answer|>
<question> $(12846+(90572/43105))\times 76230-50418$ </question>
<answer> 979566891.3525446 </answer>
<|External algorithm supervision|>
def task_calibration (ori_text):
<|Resulting Text|>
<question> $(12846+(90572/43105))\times 76230-50418$ </question>
<answer> 979360336.076325 </answer>
Example for Interviewee (Operation):
<|User Prompt|>
Please solve the given math problem and return your answer in <answer> </answer> format.
<|Interviewer’s Answer|>
<answer> 979360426.076235 </answer>
G.6 Prompt Templates for Binary Tree Traversal Tasks
Example for Interviewer:
<|System Prompt|>
You are a helpful assistant proficient in computer science algorithms.
<|User Prompt|>
Please generate a binary tree traversal problem with the following requirements:
1. The tree should have exactly {level+2} levels
2. The number of nodes should be between {2**(level+1)} and {2**(level+2)-1}(inclusive)
3. The sequences should be different from these last problems: {last problems}
4. The tree structure should be a valid binary tree
5. Each node ID should be a unique integer
Provide three types of traversal sequences. Format your response as:
<preorder> </preorder>
<inorder> </inorder>
<postorder> </postorder>
<|Interviewer’s Answer|>
<preorder> 50 30 20 10 15 13 12 40 35 70 60 80 75 78 76 79 77 </preorder>
<inorder> 12 13 15 10 20 30 35 40 50 60 70 75 76 78 77 79 80 </inorder>
<postorder> 12 13 15 10 20 35 40 30 60 76 77 79 78 75 80 70 50 </postorder>
<|External algorithm supervision|>
def task_calibration (preorder, inorder):
<|Resulting Text|>
<preorder> 50 30 20 10 15 13 12 40 35 70 60 80 75 78 76 79 77 </preorder>
<inorder> 12 13 15 10 20 30 35 40 50 60 70 75 76 78 77 79 80 </inorder>
<postorder> 12 13 15 10 20 35 40 30 60 76 77 79 78 75 80 70 50 </postorder>
Example for Interviewee:
<|User Prompt|>
Given the preorder and inorder traversals of a binary tree (with node IDs separated by spaces) <preorder> ... </preorder> and <inorder> ... </inorder>, compute its postorder traversal. Please solve the given algorithm problem and return your answer in <answer> </answer> format.
<|Interviewer’s Answer|>
<answer>12 13 15 10 20 35 40 30 60 76 77 79 78 75 80 70 50</answer>
G.7 Prompt Templates for Shortest Path Search Tasks
Example for Interviewer:
<|System Prompt|>
You are a helpful assistant proficient in computer science algorithms.
<|User Prompt|>
Please generate a shortest path search problem with the following requirements:
1. The network should be an undirected weighted graph
2. The network is given in an adjacency list format. For example, for each node, its connected nodes along with the corresponding edge weights are listed as:
<network>
A: B(2), C(4)
B: A(2), D(3)
C: A(4), D(1)
D: B(3), C(1)
</network>
3. The number of nodes is {level_node+4}, or the average edge number of each node is {level_edge+2}
4. Node IDs must be uppercase or lowercase English letters.
Please randomly select two nodes from the graph and compute the weighted shortest path between them (return the number only). Format your response as:
<Node ID1> </Node ID1>
<Node ID2> </Node ID2>
<path> </path>
<|Interviewer’s Answer|>
<network>
A: B(3), C(1), F(8)
B: A(3), D(5), E(7)
C: A(1), D(2)
D: B(5), C(2), E(1)
E: B(7), D(1), F(4)
F: A(8), E(4)
</network>
<Node ID1> A </Node ID1>
<Node ID2> E </Node ID2>
<path> 4 </path>
<|External algorithm supervision|>
def task_calibration (ori_text):
<|Resulting Text|>
<network>
A: B(3), C(1), F(8)
B: A(3), D(5), E(7)
C: A(1), D(2)
D: B(5), C(2), E(1)
E: B(7), D(1), F(4)
F: A(8), E(4)
</network>
<Node ID1> A </Node ID1>
<Node ID2> E </Node ID2>
<path> 4 </path>
Example for Interviewee:
<|User Prompt|>
Given the nodes <Node ID1> A </Node ID1> and <Node ID2> E </Node ID2> of a undirected weighted graph <network> </network>.
The network is given in an adjacency list format. For example, for each node, its connected nodes along with the corresponding edge weights are listed as:
<network>
A: B(2), C(4)
B: A(2), D(3)
C: A(4), D(1)
D: B(3), C(1)
</network>
Please calculate the weighted shortest path between them and return your answer in <answer> </answer> format with the number only.
<|Interviewer’s Answer|>
<answer> 4 </answer>
G.8 Prompt Templates for Code Generation Tasks
Example for Interviewer:
<|System Prompt|>
You are a helpful programming assistant and an expert Python programmer.
<|User Prompt|>
Please generate a common coding problem specification without any code and the test input as well as the corresponding printable output in the following format:
<problem> </problem>
<input> </input>
<output> </output>
Note that the new problem should be different from the {former problems}.
<|Interviewer’s Answer|>
<problem> Given a list of integers, determine whether any two distinct numbers in the list add up to a specific target number. Return ‘True’ if such a pair exists, otherwise return ‘False’. </problem>
<input> numbers = [2, 7, 11, 15], target = 9 </input>
<output> True </output>
Example for Interviewee:
<|User Prompt|>
Given a coding question and the test input as well as the corresponding printable output in the following format:
<problem> </problem>
<input> </input>
<output> </output>
Please generate a correct Python program that matches the specification and output requirements. The code must be directly executable, so please import the necessary library files and pay attention to code indentation and line breaks. Return the code for two parts in the following format:
<code_function> </code_function>
if __name__ == ‘__main__’:
<code_main> </code_main>
<|Interviewee’s Answer|>
<code_function>
from typing import List
def has_pair_with_sum(numbers: List[int], target: int) -> bool:
seen = set()
for num in numbers:
complement = target - num
if complement in seen:
return True
seen.add(num)
return False
</code_function>
<code_main>
numbers = [2, 7, 11, 15]
target = 9
print(has_pair_with_sum(numbers, target))
</code_main>
G.9 Prompt Templates for Code Debugging Tasks
Example for Interviewer:
<|System Prompt|>
You are a helpful programming assistant and an expert Python programmer.
<|User Prompt|>
First, please generate a common coding problem specification and the test input as well as the corresponding printable output in the following format:
<problem> </problem>
<input> </input>
<output> </output>
Then, generate a correct Python program that matches the specification and output requirements. The code must be directly executable, so please import the necessary library files and pay attention to code indentation and line breaks. Return the code for two parts in the following format:
<code_function> </code_function>
if __name__ == ‘__main__’:
<code_main> </code_main>
Note that the new problem should be different from the {former problems}.
<|Interviewer’s Answer|>
<problem> Given a list of integers, determine whether any two distinct numbers in the list add up to a specific target number. Return ‘True’ if such a pair exists, otherwise return ‘False’. </problem>
<input> numbers = [2, 7, 11, 15], target = 9 </input>
<output> True </output>
<code_function>
from typing import List
def has_pair_with_sum(numbers: List[int], target: int) -> bool:
seen = set()
for num in numbers:
complement = target - num
if complement in seen:
return True
seen.add(num)
return False
</code_function>
<code_main>
numbers = [2, 7, 11, 15]
target = 9
print(has_pair_with_sum(numbers, target))
</code_main>
<|External code masking|>
def code_masking (code, masking_ratio):
<|Resulting Content|>
<code_function>
from ty***g im**rt List
def has_pair_with_sum(numbers: List[int], target: int) -> bool:
seen = set()
for num in numbers:
comp***ent = t**get - num
if comp**ment in seen:
return True
seen.******
return False
</code_function>
Example for Interviewee:
<|User Prompt|>
Given a coding question and the test input as well as the corresponding printable output in the following format:
<problem> </problem>
<input> </input>
<output> </output>
Please repair the masked Python code <code_function> </code_function> to match the specification and output requirements. The code must be directly executable, so please import the necessary library files and pay attention to code indentation and line breaks. Then generate a fixed version of the program in the following format:
<code_function> </code_function>
<|Interviewee’s Answer|>
<code_function>
from typing import List
def has_pair_with_sum(numbers: List[int], target: int) -> bool:
seen = set()
for num in numbers:
complement = target - num
if complement in seen:
return True
seen.add(num)
return False
</code_function>